# In Search of Cosmic Time: Complete Observables and the Clock Hypothesis

### v1

@@ Abstract.
This paper consider a new and deeply challenging face of the problem of time in the context of cosmology drawing on the work of Thiemann (2006). Thiemann argues for a radical response to cosmic problem of time that requires us to modify the classical Friedmann equations. By contrast, we offer a conservative proposal for solution of the problem by bringing together ideas from the contemporary literature regarding reference frames (Bamonti 2023; Bamonti and Gomes 2024), complete observables (Gryb and Thébault 2016, 2023), and the model-based account of time measurement (Tal 2016). On our approach, we must reinterpret our criteria of observability in light of the clock hypothesis and the model-based account of measurement in order to preserve the Friedmann equations as the dynamical equations for the universe.
@@ 1. Introduction
Spacetime symmetry and time evolution are not straightforward to reconcile. Most pointedly, the spacetime diffeomorphism invariance of general relativity would appear to imply, recalling Minkowski (1908)’s famous remark, that time by itself is doomed to fade away into mere a shadow. [^1] The implications of diffeomorphism invariance and the related notions of general covariance and background independence is much debated. See Pooley (2017); James Read (2023). In the context of cosmology, however, it is standard to make reference to a notion of cosmic time. [^2] A cosmic time function on a spacetime manifold $M$ assigns to any $p\in M$ the supremum of the durations of all future-directed continuous timelike curves ending at $p$ (Andersson et al. 1998; Fletcher 2025). See Smeenk (2013); Callender and McCoy (2021) for philosophical overviews of time in cosmology. See Rugh and Zinkernagel (2009) for a discussion focused on cosmic time. One application of this time is of course in discussions of the age of the universe. Cosmic times can be shown to be well-defined in any relativistic cosmology with stable causal structure (Hawking 1969). Moreover, when they can be defined, such times manifestly are consistent with the spacetime symmetries of general relativity since they are spacetime diffeomorphism invariant. There are, however, plausible formal and physical reasons to doubt the status of a cosmic time as an independent physical observable. First, on very general grounds, we might demand of an observable that it corresponds to a measurable quantity of the theory and it is not immediately clear how one would measure cosmic time. Furthermore, we might also demand that genuine observables are relational in the sense that they correspond to correlations between measurable quantities. On such an understanding of what is to be an observable we would not expect that cosmic time is observable since it is does not appear to have the right form to be a correlation between two measurable quantities.
@@
This potential challenge to the observable status of cosmic time becomes all the more acute in the context of the canonical formalism for general relativity. A remarkable achievement of the canonical gravitational formalism, pioneered by Dirac and ADM (Dirac 1958b; Arnowitt et al. 1959, 1962), was to show that one can represent spacetime symmetries within a 3D+1 space and time formalism in constrained Hamiltonian terms, with the spacetime diffeomorphism symmetry of the 4D covariant theory represented via hyper-surface deformation algebroid of constraints. Although it is restricted to spacetimes admitting globally hyperbolic topology, there is no expectation that such a limitation will conflict with the application of the formalism to cosmology (the black hole case is more subtle due to the Kerr solutions not being globally hyperbolic).333 “There are some good reasons for believing that all physically realistic spacetimes must be globally hyperbolic (see Penrose (1980))” (Wald 1984, p. 200). A further attractive feature of the canonical formulation of general relativity is that it comes equipped with a formal criterion for a quantity to be an observable in terms of the Dirac criterion. Furthermore, in a groundbreaking and highly influential analysis it was shown that a particular approach to relational concept of an observable, as a correlation between two partial observables leading to a complete observable (Rovelli 2002b), provides a construction of objects that satisfy the Dirac criterion. An unambiguous criterion for functions to be observables is of course indispensable staging material needed towards the pursuit of canonical quantization. It is the algebra of classical observables that one seeks to faithfully represent as operators on a physical Hilbert space. Such a criterion is also useful for settling questions of observability in the classical formalism. The problem, however, is that it is not at all clear that canonical representations of cosmic time would in fact satisfy the Dirac criterion. In particular, prima facie, cosmic time does not appear to be the right kind of relational object to fit with the complete observables approach. It is not at all clear that cosmic time is either a partial or complete observable in Rovelli’s terms. This is, ultimately, a cosmological manifestation of the infamous problem of time, which we name the cosmic problem of time and upon which we will have more to say later.
@@
The acuteness of the problem of cosmic time can be illustrated most clearly in the context of the Friedmann equations that are the cornerstone of the standard model of cosmology. These equations describe the expansion of the universe in terms of the dynamics of the spatial scale factor $a$ with respect to the cosmic time parameter $t$. These equations can be straightforwardly represented in canonical terms since they turn out to simply be equivalent to the Hamilton equations. In this case it is simple to observe that neither the cosmic time nor the scale factor are Dirac observables. By the Dirac criterion, observables must be phase space functions which have (weakly) vanishing Poisson bracket with the Hamiltonian constraint that generates the dynamics of the Friedmann equation. However, clearly neither $a$ nor $t$ can satisfy such a criterion: $a$ has non-zero Poisson bracket and $t$ is not even a phase space function. Though simple to articulate and difficult to ignore once it is articulated, this problem has received almost no discussion in physics or philosophy. The notable exception is the treatment of Thiemann (2006), who explicitly argues that ‘it is incorrect to interpret the FLRW equations as evolution equations of observable quantities’ (p. 9). Rather, he suggests, we should follow the complete observables procedure and re-write the equations relationally. The true evolution equations would then acquire observable modifications when compared to the Friedmann equations. We are thus lead from a formal-conceptual problem regarding time and observability to a physical proposal for a new approach to cosmology with empirical consequences.
@@
This paper will consider this chain of argument and evaluate whether there is an alternative to Thiemann’s radical cosmological revisionism. We formulate and critically appraise a conservative approach to finding cosmic time based upon the application of the clock hypothesis in the context of the concept of Hubble flow. The development of this alternative approach will lead us, in turn, to reconsider the methodological status of the clock hypothesis in cosmology and the meaning of ‘measurability’ in the context of the complete observables programme. Our analysis draws crucially on the different types of reference frames in physical theory and the insights that this can deliver for clarifying foundational questions regarding the construction of ‘complete observables’ as correlations between ‘partial observables’. Our goal is to synthesise key ideas from the contemporary literature regarding reference frames (Bamonti 2023; Bamonti and Gomes 2024), complete observables (Gryb and Thébault 2016, 2023), and the model-based account of time measurement (Tal 2016), whilst drawing attention to a new and deeply challenging face of the problem of time in the context of cosmology. The cosmic problem of time leads to a dilemma: we can apply a conservative understanding of Dirac observables, downplay the significance of the clock hypothesis, and modify the Friedmann equations; or we can reinterpret our criteria of observability in light of the clock hypothesis and the model-based account of measurement, and preserve the Friedmann equations. Whatever option we take, something must change, for things to stay as they are.
@@
Our arguments are structured as follows. In Section 2 we provide a brief overview of the complete and partial observables programme as a response to the problem of time that seeks to preserve the Dirac criterion for observables. This will include an explicit toy model construction of a complete observable to give the reader a clear formal intuition regarding the mathematical form of these objects. We then apply recent work on reference frames to disambiguate two important details in the definition of a partial observable and better understand what it means for a physical variable to play the role of a clock in the context of a complete observable. The following Section 3 considers the status of time in the Friedmann equations, poses Thiemann’s challenge to the standard interpretation of these equations as dynamical equations and reviews his solution in terms of a de-parameterization via a phantom field through the well-known Brown and Kuchař (1995) mechanism. Section 4 then introduces the crucial idea of Hubble flow and seeks to reframe comic time as a proper time parameter $\tau$ along the Hubble flow. In this context, we consider the question of the observability and measurability of Hubble parameter (i.e. $H=\dot{a}/a$) and the question of whether we can consider $H(\tau)$ to be a complete observable whose dynamics is described by the unmodified Friedman equations. We articulate the specific problem of there being no physical system which can be understood to even approximately measure proper time along the Hubble flow. This leads, in Section 5, to the final step in the articulation of our proposed solution: the introduction of a more liberalised sense of ‘measurement procedure’ in the context of cosmic time. This more liberalised notion draws upon the model-based account of time measurement developed in the context of atomic clocks and the measurement of Coordinated Universal Time due to Tal (2016). An appendix provides a formal analysis of the status of cosmic proper time as a gauge invariant quantity and relevance of gauge fixings which supplements and supports the the analysis provided in the body of the paper.
@@ 2. Complete Observables and Real Reference Frames
The problem of time is best understood as a cluster of formal, physical and conceptual challenges to the isolation of the physical degrees of freedom in theories which display temporal diffeomorphism symmetry. [^4] See Kuchar̆ (1992); Isham (1993); Anderson (2017) for scientific overview. Casadio et al. (2024) provides an overview of a family of alternative approaches in which first-class phase-space constraints may be relaxed based on an interpretation of them as fixing the values of new degrees of freedom. Technically informed discussion in the philosophical literature include Belot and Earman (2001); Belot (2007); Gryb and Thébault (2016); Thébault (2021b). A hybrid formal and philosophical monograph length treatment of the global problem of time is Gryb and Thébault (2023). Further references will be given where relevant below. In the canonical representation, many of the challenges stem from lack of an unambiguous phase space representation of re-foliation symmetry and the implications that this has for quantization. However, the problem is not restricted to canonical representations, and reoccurs in covariant form, for example, in terms of the challenge of finding an appropriate measure in path-integral approaches. One particular pressing aspect of the problem is the tension between the standard definition of a gauge invariant observable and the seemingly obvious fact that observable quantities change. In particular, in the context of constrained Hamiltonian theories, following Dirac (1950, 1958a, 1964) the criterion to be an observable is to have (weakly) vanishing Poisson bracket with first class (primary) constraints. [^5] This idea also traces back to the discussions of Bergmann (1956); Bergmann and Komar (1960); Bergmann (1961b, a); Bergmann and Komar (1962) and so one might plausibly use the term Bergmann-Dirac observables. However, Bergmann changed his view at various points. See Pitts (2019) for discussion. Even for simple theories, temporal diffeomorphism immediately leads to a problem of time since, in such theories, we have that the Hamiltonian is a sum of first class constraints. Application of the Dirac criterion then immediately implies that observables are condemned to be frozen as constants of the motion. In general relativity this problem re-occurs in a more complex fashion but with essentially the same elements. There are an infinite family of Hamiltonian constraints and if we insist that Dirac observables commute with them, then the observables of general relativity are frozen.
@@
The standard approach to the problem of reconciling the Dirac definition of observables with the necessity to describe dynamics is the partial and complete observables approach. This approach was pioneered by Rovelli (1991a, b, 2002b, 2004, 2007) and later formalized by Dittrich (2006, 2007). [^6] For detailed overview see (Thiemann 2007; Tambornino 2012). Important developments of the approach include (Gambini and Porto 2001; Gambini et al. 2009) and (Bojowald et al. 2011a, b; Höhn 2019). Critical responses include (Kuchar̆ 1991, 1992; Kuchar 1999; Dittrich et al. 2017). For a review of the various notions of observable, that includes discussion of the limitations of the partial and complete observables approach, see (Anderson et al. 2014; Anderson 2017). For philosophical analysis of the ontological implications of the partial and complete observables approach an excellent extended discussion can be found in (Rickles 2007, pp. 161–171). For a further overview see (Thébault 2021b), which contains further references and discussion. The essence of such approach is to designate a subset of measurable quantities or ‘partial observables’ as internal clocks, and then use these clocks to construct ‘complete observables’ that are both predicable and measurable, and which correspond to Dirac observables. More specifically, the formal application of the approach requires one to consider, for each Hamiltonian constraint, one physical variable to play the role of a physical clock. One first constructs expressions for the clock and non-clock variables in terms of the parameterized flow of the constraints; the simplest way to do this is via the relevant Hamilton-Jacobi equation (Rovelli 2004; Gryb and Thébault 2016). One next inverts the flow equation for the clock variable and substitutes it into flow equations for the other variables to construct an algebraic expression for their correlation that is parameter free. Finally, one considers the correlation between the clock variable and the other variables at a particular value of the clock variable. This is a complete observable and corresponds to the value of the non-clock partial observables when the clock partial observables takes a particular value.
@@
Let us provide the simplest possible physical example so the reader can conceptualise clearly how the procedure works. Consider two free particles moving in one dimension and described by a theory with a single Hamiltonian constraint. We can write an expression for the integral of motion in terms of the time parameter of the flow of the constraint, by solving the Hamilton-Jacobi equation. This will give us an expression for the position of each particle, $q_{i}$, as a functions of the constants of motion (i.e. inital position and momenta), $Q_{i}$ and $P_{i}$, and parameter time, $t$. This takes the form:
@@
for $i=1,2$. These variables are partial observables and do not commute with the Hamiltonian constraint $\mathcal{H}$, since we have that $\{\mathcal{H},q_{i}\}=\dot{q}_{i}\neq 0$. However, we can combine the two expressions to describe the correlation between the values of the position of each particle. We do this by inverting the expression for one variable such that we obtain $t$ as a function of $(q_{i},Q_{i},P_{i})$, and then inserting this expression into the expression for the other. The first variable is then playing the role of a physical clock and we evaluate the second variable for a given value of the second variable, say $s\in\mathbb{R}$. In this way we get a family of complete observables, one for each value of $s$ [^7] Note that this is also the construction of the so-called evolving constant of motions (Rovelli 1991a). In fact, a complete observable formally coincides with an evolving constant. The difference between the two concepts lies mainly in the fact that for evolving constants, the focus is on the evolution of the quantity with respect to the parameter $s$ that serves as ‘internal time’.
For even slightly complicated physical systems the inversion step may run into significant obstacles and is typically such that we can only define the relevant expressions for restricted values of the time parameter. Dittrich (2007) provides a detailed treatment of such a case. In our case, by contrast, since the physical dynamics is trivial and we are able to solve Hamilton equations for the considered system, the inversion is simply given by:
@@
Finally, we evaluate our expression $q_{2}(q_{1})$ at $q_{1}=s$ to get the parameterized family of complete observables:
@@
There is a specific tension within the physics literature regarding the interpretation of the partial observables. This tension will prove crucial in the context of application of the approach to cosmology. Consider, in particular, that according to the original approach of Rovelli (2002b), by definition, a partial observable is ‘a physical quantity with which we can associate a (measuring) procedure leading to a number’ (p. 2). By contrast, following Thiemann (2007) we have that ‘a measurable quantity is always a complete observable, even pointers of a clock are observables and not partial observables. Now complete observables are defined with respect to non-measurable quantities…which we will simply call non-observables’ (p. 78). A third view is advocated by Gryb and Thébault (2016, 2023) in the context of theories with a single Hamiltonian constraint. On this approach, one can think of the complete and partial observables programme as allowing us de-parameterize evolution purely in terms of observable quantities. However, this evolution is fundamentally controlled by the evolution equations generated by the Hamiltonian constraint and is always well-defined, even when a particular deparameterization breaks down. On this approach, even if one wishes to use parameter-free complete observable expressions, one is still required to retain the full partial observables representation. This supports the Rovelli (2002b) perspective, in which partial observables are measurable quantities, rather than the Thiemann (2007) perspective, where the partial observables are understood as non-measurable.
A further important disambiguation can be made based upon the connection between partial and complete observables and reference frames. The role of reference frames in general relativity has an extensive philosophical literature. [^8] See Earman and Friedman (1973); Earman (1974), Norton (1989, 1993), DiSalle (2020). Recently, a community of scholars has also emerged in the field of the so-called quantum reference frames, see Giacomini (2021); Kabel et al. (2024) and reference therein. Most relevant to our analysis is the distinction made by Bamonti (2023) between: ‘Idealised Reference Frames’ (IRFs), in which any dynamical interaction of the material system represented by the reference frame is ignored; ‘Dynamical Reference Frames’ (DRFs), in which the set of equations that determine the dynamics of the matter field is included but the the stress-energy tensor of the matter field used as reference frame is neglected; and ‘Real Reference Frames’ (RRFs) in which both the dynamics of the chosen material system and its stress-energy tensor are taken into account.

### v2

@@ Abstract.
This paper considers a new and deeply challenging face of the problem of time in the context of cosmology drawing on the work of Thiemann (2006, 2007). Thiemann argues for a radical response to the cosmic problem of time that requires us to modify the classical Friedmann equations. By contrast, we offer a conservative proposal for solution of the problem by bringing together ideas from the contemporary literature regarding reference frames (Bamonti 2023; Bamonti and Gomes 2024), complete observables (Gryb and Thébault 2016b, 2023), and the model-based account of time measurement (Tal 2016). On our approach, we must reinterpret our criteria of observability in light of the clock hypothesis and the model-based account of measurement in order to preserve the Friedmann equations as the dynamical equations for the universe.
@@ 1. Introduction
1.1. Précis
  • A Cosmic Problem: In every causally stable spacetime of general relativity there exists a symmetry invariant cosmic time function. However, it is unclear whether we should treat such cosmic times as ‘observables’ according to the standard ‘Dirac criterion’ of observability deployed by scientists in analysing the theory.
  • A Radical Proposal: Taking the problem of the observability of cosmic time seriously in the context of the Friedmann equations for cosmology leads to potentially significant modifications to the equations. In particular, one can show that consistently applying the standard Dirac criterion may modify the form of the equations in a manner that can have empirical consequences.
  • Our Alternative Approach: A different option is to reinterpret the criteria of observability in light of the clock hypothesis. On the assumption that there are clocks which measure proper time along any given world-line, one is able to treat cosmic time as an observable and preserve the form of the Friedmann equations.
  • The Conservative’s Dilemma: Whatever option we take, something must change, for things to stay as they are.
1.2. The Problem of Cosmic Time
Spacetime symmetry and time evolution are not straightforward to reconcile in the context of the diffeomorphism invariance of general relativity. In particular, the ability to ‘re-slice’ relativistic spacetimes into arbitrary decompositions of space-like hypersurfaces indicates that time is ‘many-fingered’ within the theory. [^1] The implications of diffeomorphism invariance and the related notions of general covariance and background independence are much debated. See Pooley (2017); Read (2023). Many-fingered time and the decomposition into space-like hyperspaces requires relativistic spacetimes to be globally hyperbolic (Geroch 1970; Misner et al. 2017). In the context of cosmology, however, it is standard to make reference to a notion of ‘cosmic time’. One application of cosmic time is, of course, in discussions of the age of the universe. Most vividly, in simple FLRW cosmological models the cosmic time function labels isotropic and homogenous spatial hypersurfaces and corresponds to the temporal length of geodesic paths that run from the big bang until the present day. [^2] An important connected notion of time in general relativity is ‘York time’. This corresponds to the time that labels spatial slices of constant mean curvature and has important connections with conformal invariant approaches to the initial value problem of general relativity (York 1972, 1973; Gomes et al. 2011; Gryb and Thébault 2016c). York time corresponds to cosmic time for homogenous spatial slices, meaning that in FLRW models the constant mean curvature slices are the same as the slices of constant cosmic time. However, in general, the two are distinct (see Roser and Valentini (2014)). Cosmic times can be shown to be well-defined in any relativistic cosmology with stable causal structure (Hawking 1969) and are consistent with the spacetime symmetries of general relativity since they can be defined in a manifestly spacetime diffeomorphism invariant manner. [^3] Formally, a cosmic time function on a spacetime manifold $M$ is a function that assigns to any $p\in M$ the supremum of the durations of all future-directed continuous timelike curves ending at $p$ (Andersson et al. 1998; Fletcher 2025). See Smeenk (2013); Callender and McCoy (2021) for philosophical overviews of time in cosmology. See Rugh and Zinkernagel (2009) for a discussion focused on cosmic time.
There are, however, plausible formal and physical reasons to doubt the status of a cosmic time as an independent physical observable. First, on very general grounds, we might demand of an observable that it corresponds to a measurable quantity of the theory and it is not immediately clear how one would measure cosmic time since there is no natural ‘clock’ that we would understand to read off cosmic time nor clear methodology to construct one. Second, we might also demand that genuine observables are relational in the sense that they correspond to correlations between measurable quantities. On such an understanding of what is to be an observable we would not expect that cosmic time is observable since it is does not appear to have the right form to be a correlation between two measurable quantities.
@@ 1.2. The Problem of Cosmic Time
This potential challenge to the observable status of cosmic time becomes all the more acute in the context of the canonical formalism for general relativity. A remarkable achievement of the canonical gravitational formalism, pioneered by Dirac and ADM (Dirac 1958b; Arnowitt et al. 1959, 1962), was to show that one can represent spacetime symmetries within a 3D+1 space and time formalism in constrained Hamiltonian terms. In this formulation, the spacetime diffeomorphism symmetry of the 4D covariant theory is encoded in the hypersurface deformation algebroid of constraints. This algebroid is a mathematical structure encoding how the constraints governing the dynamics of the 3D spatial slices transform under infinitesimal normal and tangential deformations of the slices (Teitelboim 1973; Gryb and Thébault 2016a). These deformations correspond to ‘shifts’ in time and space and ensure that the canonical Dirac formalism preserves the underlying covariance of general relativity, even within the 3D+1 decomposition. Although it is restricted to spacetimes admitting globally hyperbolic topology, there is no expectation that such a limitation will conflict with the application of the formalism to cosmology (the black hole case is more subtle due to the Kerr solutions not being globally hyperbolic). [^4] The general expectation that all physically realistic spacetimes must be globally hyperbolic is based upon assumptions regarding the instability of features such as Cauchy horizons that lead to a breakdown in global hyperbolicity. See Penrose 1980, §12, (Wald 1984, p. 202). A further attractive feature of the canonical formulation of general relativity is that it comes equipped with a formal criterion, known as the Dirac criterion, which identifies a quantity as an observable if it has (weakly) vanishing Poisson bracket with all first-class constraints of the theory. Furthermore, in a groundbreaking and highly influential analysis, it was shown that a particular approach to relational conception of an observable, as a correlation between two partial observables leading to a complete observable (Rovelli 2002b), provides a construction of objects that satisfy the Dirac criterion. An unambiguous criterion for functions to be observables is of course indispensable staging material needed towards the pursuit of canonical quantisation. It is the algebra of classical observables that one seeks to faithfully represent as operators on a physical Hilbert space. Such a criterion is also useful for settling questions of observability in the classical formalism. The problem, however, is that it is not at all clear that canonical representations of cosmic time would in fact satisfy the Dirac criterion. In particular, prima facie, cosmic time does not appear to be the right kind of relational object to fit with the complete observables approach. It is not at all clear that cosmic time is either a partial or complete observable in Rovelli’s terms. This is, ultimately, a cosmological manifestation of the infamous problem of time, which we name the cosmic problem of time and upon which we will have more to say later.
@@
The acuteness of the cosmic problem of time can be illustrated most clearly in the context of the Friedmann equations that are the cornerstone of the standard model of cosmology. These equations describe the expansion of the universe in terms of the dynamics of the spatial scale factor, $a$, with respect to the cosmic time parameter, $t$. These equations can be straightforwardly represented in canonical terms since they turn out to simply be equivalent to the Hamilton equations. In this case it is simple to observe that neither the cosmic time nor the scale factor are Dirac observables. As per our discussion above, by the Dirac criterion, such observables must be phase space functions which have (weakly) vanishing Poisson bracket with the Hamiltonian constraint that generates the dynamics of the Friedmann equation. However, clearly neither $a$ nor $t$ can satisfy such a criterion. This is because, in the first case, $a$ evidently has non-zero Poisson bracket with the Hamiltonian for any non-trivial cosmological dynamics by the definition of the Friedmann equations. In the second case, not only is $t$ not naturally understood as a phase space variable, rather it is a parameter of a phase space flow, but any phase space function that ‘marches in step’ with $t$ will have non-zero Lie drag with the Hamilton vector field of the Hamiltonian, and thus non-vanishing Poisson bracket. Though simple to articulate and difficult to ignore once it is recognised, this problem has received almost no discussion in physics or philosophy.
Isolation of the problem crucially depends upon how strict one is regarding the application of the Dirac criterion in the context of ‘partial observables’ – both in general and in the cosmological case. Important discussions of this issue in general and in the cosmological context can be found in the work of the noted theoretical physicist Thomas Thiemann. The general question of observability of partial observables is discussed in his textbook Thiemann (2007). There he argues that ‘a measurable quantity is always a complete observable, even pointers of a clock are observables and not partial observables. Now complete observables are defined with respect to non-measurable quantities…which we will simply call non-observables’ (p. 78). Thiemann’s view implies that both the scale factor and the cosmic time parameter are non-observables and that the Friedman equations therefore do not describe the evolution of observable quantities. Indeed, an analysis along precisely these lines can be found in Thiemann (2006). There Thiemann explicitly argues that ‘it is incorrect to interpret the FLRW equations as evolution equations of observable quantities’ (p. 9). Rather, he suggests, we should follow the complete observables procedure and re-write the equations relationally. The true evolution equations can then shown for a simple model to acquire empirical modifications when compared to the Friedmann equations. We are thus lead from a formal-conceptual problem regarding time and observability to a physical proposal for a new approach to cosmology with empirical consequences. One of the major goals of this paper is to consider this chain of argument leading to Thiemann’s radical cosmological revisionism.
@@
The second major goal of this paper is to formulate and defend a conservative approach to finding cosmic time based upon the application of the clock hypothesis in the context of the concept of Hubble flow. [^5] The idea of ‘operationalising’ cosmic time via Hubble flow is a well-known standard approach in the literature on FLRW cosmology. See e.g. (Misner et al. 2017, § 27.4) or also Ryden (2016). The development of this alternative approach will lead us, in turn, to reconsider the methodological status of the clock hypothesis in cosmology and the meaning of ‘measurability’ in the context of the complete observables programme. Our analysis draws crucially on the different types of reference frames in physical theory and the insights that this can deliver for clarifying foundational questions regarding the construction of ‘complete observables’ as correlations between ‘partial observables’. Our goal is to synthesise key ideas from the contemporary literature regarding reference frames (Bamonti 2023; Bamonti and Gomes 2024), complete observables (Gryb and Thébault 2016b, 2023), and the model-based account of time measurement (Tal 2016), whilst drawing attention to a new and deeply challenging face of the problem of time in the context of cosmology. The cosmic problem of time leads to a dilemma: we can apply a conservative understanding of Dirac observables, downplay the significance of the clock hypothesis, and modify the Friedmann equations; or we can reinterpret our criteria of observability in light of the clock hypothesis and the model-based account of measurement, and preserve the Friedmann equations. Whatever option we take, something must change, for things to stay as they are.
1.3. Roadmap
@@
Section 2 provides a brief overview of the complete and partial observables programme as a response to the problem of time. We then apply recent work on reference frames to disambiguate two important details in the definition of a partial observable and better understand what it means for a physical variable to play the role of a clock in the context of a complete observable. Section 3 considers the status of time in the Friedmann equations, poses Thiemann’s challenge to the standard interpretation of these equations as dynamical equations and reviews his solution in terms of a de-parametrisation via a phantom field through the well-known Brown and Kuchař (1995) mechanism. Section 4 reframes comic time as a proper time parameter $\tau$ along the Hubble flow. In this context, we consider the question of the observability and measurability of Hubble parameter (i.e. $H=\dot{a}/a$) and the question of whether we can consider $H(\tau)$ to be a complete observable whose dynamics is described by the unmodified Friedman equations. Problematically, there is no physical system which even approximately measures proper time along the Hubble flow. Section 5 articulates our proposed solution: the introduction of a more liberalised sense of ‘measurement procedure’ in the context of cosmic time. This more liberalised notion draws upon the model-based account of time measurement developed in the context of atomic clocks by Tal (2016).
@@ 2. Complete Observables and Real Reference Frames
The problem of time is best understood as a cluster of formal, physical and conceptual challenges to the isolation of the physical degrees of freedom in theories which display temporal diffeomorphism symmetry. [^6] See Kuchar̆ (1992); Isham (1993); Anderson (2017) for scientific overview. Casadio et al. (2024) provides an overview of a family of alternative approaches in which first-class phase-space constraints may be relaxed based on an interpretation of them as fixing the values of new degrees of freedom. Technically informed discussion in the philosophical literature include Belot and Earman (2001); Belot (2007); Thébault (2021b). A hybrid formal and philosophical monograph-length treatment of the global problem of time is Gryb and Thébault (2023). Further references will be given where relevant below. In the canonical representation, many of the challenges stem from lack of an unambiguous phase space representation of re-foliation symmetry and the implications that this has for quantisation. However, the problem is not restricted to canonical representations, and reoccurs in covariant form, for example, in terms of the challenge of finding an appropriate measure in path-integral approaches. One particular pressing aspect of the problem is the tension between the standard definition of a gauge-invariant observable and the seemingly obvious fact that observable quantities change. In particular, in the context of constrained Hamiltonian theories, following Dirac (1950, 1958a, 1964) the criterion to be an observable is to have (weakly) vanishing Poisson bracket with first class (primary) constraints. [^7] This idea also traces back to the discussions of Bergmann (1956); Bergmann and Komar (1960); Bergmann (1961b, a); Bergmann and Komar (1962) and so one might plausibly use the term Bergmann-Dirac observables. However, Bergmann changed his view at various points. See Pitts (2019) for discussion. Even for simple theories, temporal diffeomorphism symmetry immediately leads to a problem of time since, in such theories, we have that the Hamiltonian is a sum of first class constraints. Application of the Dirac criterion then immediately implies that observables are condemned to be frozen as constants of the motion. In general relativity this problem recurs in a more complex fashion but with essentially the same elements. There is an infinite family of Hamiltonian constraints and if we insist that Dirac observables commute with them, then the observables of general relativity are frozen.
@@
The standard approach to the problem of reconciling the Dirac definition of observables with the necessity to describe dynamics is the partial and complete observables approach. This approach was pioneered by Rovelli (1991a, b, 2002b, 2004, 2007) and later formalised by Dittrich (2006, 2007). [^8] For detailed overview see (Thiemann 2007; Tambornino 2012). Important developments of the approach include (Gambini and Porto 2001; Gambini et al. 2009) and (Bojowald et al. 2011a, b; Höhn 2019). Critical responses include (Kuchar̆ 1991, 1992; Kuchar 1999; Dittrich et al. 2017). For a review of the various notions of observable, that includes discussion of the limitations of the partial and complete observables approach, see (Anderson et al. 2014; Anderson 2017). For analysis of relationship between the partial and complete observables approach and earlier influential work by Page and Wootters (1983) see the analysis of Höhn et al. (2021) – this work demonstrates the equivalence between the two approaches by putting the latter on a more rigours footing and resolving problems within its presentation as highlighted by Kuchar̆ (1992). For philosophical analysis of the ontological implications of the partial and complete observables approach an excellent extended discussion can be found in (Rickles 2007, pp. 161–171). For a further overview see (Thébault 2021b), which contains further references and discussion. The essence of this approach is to designate a subset of measurable quantities or ‘partial observables’ as internal clocks, and then use correlations between the remaining variables and these clocks to construct ‘complete observables’ that are both predicable and measurable, and which correspond to Dirac observables. More specifically, the formal application of the approach requires one to consider, for each Hamiltonian constraint, one physical variable to play the role of a physical clock. One first constructs parametrised flow expressions for the ‘evolution’ of the clock and non-clock variables under the Lie flow of the vector field associated with the constraint; the simplest way to do this is via the relevant Hamilton-Jacobi equation (Rovelli 2004; Gryb and Thébault 2016b, 2023). One next inverts the flow equation for the clock variable and substitutes it into flow equations for non-clock variables to construct an algebraic expression for their correlation that is parameter free. Finally, one considers the correlation between the clock variable and the other variables at a particular value of the clock variable. This is a complete observable and corresponds to the value of the non-clock partial observables when the clock partial observables takes a particular value.
@@
Let us provide the simplest possible physical example so the reader can conceptualise clearly how the procedure works. Consider two free particles moving in one dimension and described by a theory with a single Hamiltonian constraint. We can write an expression for the integral of motion in terms of the time parameter of the flow of the vector field associated with constraint, by solving the Hamilton-Jacobi equation. This will give us an expression for the position of each particle, $q_{i}$, as a functions of the constants of motion (i.e. initial position and momenta), $Q_{i}$ and $P_{i}$, and parameter time, $t$. This takes the form:
@@
for $i=1,2$. These variables are partial observables and do not commute with the Hamiltonian constraint $\mathcal{H}$, since we have that $\{\mathcal{H},q_{i}\}=\dot{q}_{i}\neq 0$. However, we can combine the two expressions for for $i=1,2$ to describe the correlation between the values of the position of each particle. We do this by inverting the expression for one variable such that we obtain $t$ as a function of $(q_{i},Q_{i},P_{i})$, and then inserting this expression into the expression for the other. The first variable is then playing the role of a physical clock and we evaluate the second variable for a given value of the second variable, say $s\in\mathbb{R}$. In this way we get a family of complete observables, one for each value of $s$ [^9] Note that this is also the construction of the so-called evolving constant of motions (Rovelli 1991a). In fact, a complete observable formally coincides with an evolving constant. The difference between the two concepts lies mainly in the fact that for evolving constants, the focus is on the evolution of the quantity with respect to the parameter $s$ that serves as ‘internal time’. For even slightly complicated physical systems the inversion step may run into significant obstacles and is typically such that we can only define the relevant expressions for restricted values of the time parameter. Dittrich (2007) provides a detailed treatment of such a case. In our case, by contrast, since the physical dynamics is trivial and we are able to solve Hamilton equations for the considered system, the inversion is simply given by:
@@
Finally, we evaluate our expression $q_{2}(q_{1})$ at $q_{1}=s$ to get the parametrised family of complete observables:
@@
There is a specific tension within the physics literature regarding the interpretation of the partial observables. This tension will prove crucial in the context of application of the approach to cosmology. Consider, in particular, that according to the original approach of Rovelli (2002b), by definition, a partial observable is ‘a physical quantity with which we can associate a (measuring) procedure leading to a number’ (p. 2). [^10] In his original definition Rovelli makes clear that the definition should not be understood in operationalist terms. In particular, he notes ‘The operational tone of the [partial observable] definition does not imply any adherence to operationalism here (Bridgman 1927): the reference to measuring procedures is just instrumental for clarifying a distinction.’ (ibid. p. 2). Crucially, the partial observables are well-defined theoretical quantities whose definition does not require specification of a measurement procedure. For more discussion of operationalism see Chang (2021); Fankhauser and Dürr (2021). By contrast, following Thiemann (2007) we have that ‘a measurable quantity is always a complete observable, even pointers of a clock are observables and not partial observables. Now complete observables are defined with respect to non-measurable quantities…which we will simply call non-observables’ (p. 78). A third view is advocated by Gryb and Thébault (2016b, 2023) in the context of theories with a single Hamiltonian constraint. On this approach, one can think of the complete and partial observables programme as allowing us de-parametrise evolution purely in terms of observable quantities. However, this evolution is fundamentally controlled by the evolution equations generated by the Hamiltonian constraint and is always well-defined, even when a particular deparametrisation breaks down. On this approach, even if one wishes to use parameter-free complete observable expressions, one is still required to retain the full partial observables representation. This supports the Rovelli (2002b) perspective, in which partial observables are measurable quantities, rather than the Thiemann (2007) perspective, where the partial observables are understood as non-measurable.